there is a 4.8 percent coupon bond with six years to maturity and a current price of $1,048.30. What is the dollar value of an 01 for the bond? (Do not round intermediate calculations. Round your answer to 3 decimal places. Omit the "$" sign in your response.) Dollar value $
Answer #1
                  K = N
Bond Price =∑ [( Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =6
1048.3 =∑ [(4.8*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^6
                   k=1
YTM% = 3.88

Period Cash Flow Discounting factor PV Cash Flow Duration Calc
0 ($1,048.30) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period
1                 48.00                                                             1.04                    46.21                  46.21
2                 48.00                                                             1.08                    44.48                  88.96
3                 48.00                                                             1.12                    42.82                128.46
4                 48.00                                                             1.16                    41.22                164.88
5                 48.00                                                             1.21                    39.68                198.40
6           1,048.00                                                             1.26                  834.01              5,004.04
      Total              5,630.96
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=5630.96/(1048.3*1)
=5.371513
Modified duration = Macaulay duration/(1+YTM)
=5.37/(1+0.0388)
=5.170883
dollar value = -modified duration*current price*1 percent change in YTM
Dollar value = -5.17*1048.3*0.01
Dollar value = -54.2
DV01 = Dollar value*0.01
=-54.19711*0.01
-0.5419711